Grade 11Physical Chemistry

If the velocity of electron in ftrst orbit of H atom is 2.18 x10^6 m/s,then the velocity of an electron in 4th orbit of He+ ion is c =velocity of light
1)c/275 2)c/312 3)c/137 40 c/217

Ankita Roy
7 Years agoGrade 11

1 Answer

Shweta Ahire
7 Years ago
Velocity of electron = 2.18 × 10^6 × Z/n   m/sec
Where Z = atomic no. and n = orbit no. 
So velocity of electron in 4th orbit of he+ ion = 2.18 × 10^6 × 2/4 = 1.09 × 10^6 = (3 × 10^8)/275
And 3 × 10^8 =C     so ans is C/275